solve for y 3x+2y+1=4
Answer by smallfry(7)

solve for y 3x+2y+1=4
first get rid of the one
3x+2y+1=4
-1 -1
3x+2y=4-1
know simple 4-1 to 3
3x-2y=3
now get rid of x
3x-2y=3
-3x -3x
-2y=3-3x
to get y alone divid by -2
-2y=3-3x
_________
-2 -2
y=-3/2+3/2x
*** sub. turns to add because of the 2 negs.